Water flows through a venturi meter. At the constricted section where the area is 24 cm2

, the 

pressure is 10.2 N/cm2

, and at the section where the area is 64 cm2

, the pressure is 18.0 N/cm2

Determine the velocities of water in the larger and smaller pipes and the rate of flow. 


1
Expert's answer
2021-05-13T16:22:53-0400

Area at section, "A_1=64\space cm^2"

Pressure at section, "P_1=18\space N/cm^2"

Area at constricted section, "A_2=24\space cm^2"

Pressure at constricted section, "P_1=10.2\space N/cm^2"

Density of mercury, "\rho=13.6\space g/cm^3"

Flow rate, "q=\dfrac{A_1A_2}{\sqrt{A_1^2-A_2^2}}\space \sqrt{\dfrac{2(P_1-P_2)}{\rho}}"

"q=\dfrac{64\times24}{\sqrt{64^2-24^2}}\space \sqrt{\dfrac{2(18-10.2)}{13.6}}"

"q=27.72\space cm^3/s"


Velocity at section, "v_1=\dfrac{q}{A_1}=\dfrac{27.72}{64}=0.43\space cm/s"

Velocity at constricted section, "v_2=\dfrac{q}{A_2}=\dfrac{27.72}{24}=1.155\space cm/s"